OpenStudy (anonymous):

Integrate 1/(2x+3) I know the answer but i don't see how to get it? Is it a double U substitution or is it some identify I'm suppose to know Please explain step by step Thanks

OpenStudy (cwrw238):

use standard form integral of 1/ x = ln x x is replaced by a linear function 2x + 3 the rule is to multiply by 1/coefficient of x in this case 1/2 and multiply this by ln ( 2x + 3) so integral of 1/ 2x + 3 = 1/2* ln (2x + 3) + C

OpenStudy (anonymous):

You are correct, how did u know to use 1/x though? I would of never seen that

OpenStudy (cwrw238):

1 / (2x+3) is 1/x with the x replaced by (2x+3) functions of the type 1/(ax +b) can all be integrated this way you need to have this knowledge of standard form

OpenStudy (anonymous):

notice how the integration "formula" is written as \(\Large \int \frac{1}{u}du=lnu + C \) where u is a function of x ?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

it's knowing the above I mentioned, AND doing a substitution.....

OpenStudy (anonymous):

let u=2x+3 so du=2dx ---> \(\frac {1}{2} \)du = dx \(\Large \int \frac{1}{2x+3}dx = \int \frac{1}{u} \cdot (\frac{1}{2}du)=\frac{1}{2} \int \frac{1}{u}du \)